OpenStudy (anonymous):

list 38%, 8/25, and 0.41 from least to greatest

TheSmartOne (thesmartone):

\[\huge \color{blue}{Welcome~To~OpenStudy!!!}\] It looks like you are new here. This tutorial can help you get started. http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8

TheSmartOne (thesmartone):

38%=.38 8/25=32/100=.32 and .41=.41 So .32, .38 < .41

TheSmartOne (thesmartone):

So now plgging back what we actually have it would be 8/25 < 38% < .41

TheSmartOne (thesmartone):

@Mooniris